OpenStudy (anonymous):

find the general solution of the following: 1. dy/dt=t/(ye^(y+t^2)) 2. dy/dx=cosxe^(y+sinx) 3. dy/dx=2x*sqrt(1-y^2)

OpenStudy (anonymous):

Each of these equations are separable: \[\begin{array}{c} \dfrac{dy}{dt}=\dfrac{t}{e^{y+t^2}}~~\iff~~e^y\,dy=te^{-t^2}\,dt\\\\ \hline\\ \dfrac{dy}{dx}=\cos x~e^{y+\sin x}~~\iff~~ e^{-y}\,dy=\cos x~e^{\sin x}\,dx\\\\ \hline\\ \dfrac{dy}{dx}=2x\sqrt{1-y^2}~~\iff~~ \dfrac{dy}{\sqrt{1-y^2}}=2x\,dx \end{array}\]

OpenStudy (anonymous):

forrezzz